# The solving degrees for computing Gröbner bases of affine semi-regular polynomial sequences

### v2

@@ 2.3 Extension of the notion of degree of regularity
Put $d:=\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ and assume for a contradiction that $d<D:=d_{\rm reg}(\langle F^{\rm top}\rangle)$. By Lemma 4.1.4 below, we have ${\rm LM}(G_{\rm hom})={\rm LM}(G_{\rm top})_{\leq d}\subset R$. Then $\langle{\rm LM}(G_{\rm hom})\rangle_{R}$ contains ${\rm LM}(G_{\rm top})$. This can be shown as follows: For any $t\in{\rm LM}(G_{\rm top})$, there are polynomials $a_{1},\ldots,a_{m}$ in $R$ such that $t={\rm LM}(\sum_{i=1}^{m}a_{i}f_{i}^{\rm top})$. In this case, it can be easily checked that $t={\rm LM}(\sum_{i=1}^{m}a_{i}f_{i}^{h})$, and thus $t$ is divisible by some element of ${\rm LM}(G_{\rm hom})$.
Thus, as $R/\langle F^{\rm top}\rangle=R/\langle G_{\rm top}\rangle$ is Artinian, it follows from $\langle{\rm LM}(G_{\rm hom})\rangle_{R}\supset{\rm LM}(G_{\rm top})$ that there is an element $g$ in $G_{\rm hom}$ with ${\rm LM}(g)=x_{n}^{d^{\prime}}$ for some $d^{\prime}\leq d$. Then, for any monomial $t\in R_{d^{\prime}}$, as $t\succeq x_{n}^{d^{\prime}}$, it belongs to $\langle{\rm LM}(\langle F^{h}\rangle)\rangle_{R^{\prime}}(=\langle{\rm LM}(G_{\rm hom})\rangle_{R^{\prime}}=\langle{\rm LM}(G_{\rm top})_{\leq d}\rangle_{R^{\prime}})$ by its weakly reverse lexicographicness. This implies that $t$ also belongs to $\langle{\rm LM}(G_{\rm top})\rangle_{R}$, and hence $R_{d}/\langle F^{\rm top}\rangle_{d}=0$. Thus, we have $d\geq d^{\prime}\geq d_{\rm reg}(\langle F^{\rm top}\rangle)$, a contradiction.
By the same argument as above, we can show that $\mathrm{max.GB.deg}_{\prec}(F^{\rm top})=d_{\rm reg}(\langle F^{\rm top}\rangle)$ as follows. It is obvious that $\mathrm{max.GB.deg}_{\prec}(F^{\rm top})\leq d_{\rm reg}(\langle F^{\rm top}\rangle)$. Thus, we assume to the contrary that $d:=\mathrm{max.GB.deg}_{\prec}(F^{\rm top})<D:=d_{\rm reg}(\langle F^{\rm top}\rangle)$. Then, it follows from Lemma 4.1.4 below that ${\rm LM}(G_{\rm top})={\rm LM}(G_{\rm hom})_{\leq d}$ and the ideal $\langle{\rm LM}(G_{\rm top})\rangle_{R}$ has the weak reverse lexicographicness up to $d$. Since there is an element $g$ in $G_{\rm top}$ with ${\rm LM}(g)=x_{n}^{d^{\prime}}$ for some $d^{\prime}\leq d$, any monomial in $R_{d^{\prime}}$ belongs to $\langle{\rm LM}(G_{\rm top})\rangle$ and so $R_{d^{\prime}}/\langle F^{\rm top}\rangle_{d^{\prime}}=0$, which implies $d\geq d^{\prime}\geq D$, a contradiction. ∎
@@
In Lemma 2.3.5, when $R/\langle F^{\rm top}\rangle$ is Artinian, we can easily prove the equality $\mathrm{max.GB.deg}_{\prec}(F^{\rm top})=d_{\rm reg}(\langle F^{\rm top}\rangle)$ if $\langle\mathrm{LM}(\langle F^{\rm top}\rangle)\rangle$ is weakly reverse lexicographic, not assuming that $\bm{F}^{\rm top}$ is cryptographic semi-regular nor that $\langle\mathrm{LM}(\langle F^{h}\rangle)\rangle$ is weakly reverse lexicographic.
@@ 3.1 Bounded regularity of homogenized sequences
Here we revisit the main results in [30, Section 4]. For the readability, we remain the proofs. Also, as additional remarks, we explicitly give an important property of the Hilbert-Poincaré series of $R^{\prime}/\langle F^{h}\rangle$ with $R^{\prime}=R[y]$, and also give an alternative proof for [30, Theoem 7] (Theorem 3.1.1 below).
@@
Theorem 3.1.1 (Theorem 1 (1); [30, Theoem 7]).
@@
${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)={\rm HF}_{R/\langle F^{\rm top}\rangle}(d)+{\rm HF}_{{\color[rgb]{0,0,0}R^{\prime}}/\langle F^{h}\rangle}(d-1),$ (3.1.1)
@@
$K_{i}\otimes_{R^{\prime}}R\cong\bigoplus_{1\leq j_{1}<\cdots<j_{i}\leq m}R(-d_{j_{1}\cdots j_{i}})\mathbf{e}_{j_{1}\cdots{\color[rgb]{0,0,0}{j_{i}}}},$
@@
by the dimension theorem. Since $H_{0}(K_{\bullet})=R^{\prime}/\langle F^{h}\rangle$ and ${\color[rgb]{0,0,0}H_{0}(K_{\bullet}\otimes_{R^{\prime}}R)}=R/\langle F^{\rm top}\rangle$, we have the equality (3.1.1), as desired. ∎
Remark 3.1.2 (Theorem 1 (2), (3); [30, Remark 6]).
Note that, in the proof of Theorem 3.1.1, the multiplication map $H_{0}(K_{\bullet})_{d-1}\to H_{0}(K_{\bullet})_{d}$ by $y$ is injective for all $d<D$, whence ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)$ is monotonically increasing for $d<D$. On the other hand, since $H_{0}(K_{\bullet}\otimes_{R^{\prime}}R)_{d}=(R/\langle F^{\rm top}\rangle)_{d}=0$ for all $d\geq D$ by the definition of the degree of regularity, the multiplication map $H_{0}(K_{\bullet})_{d-1}\to H_{0}(K_{\bullet})_{d}$ by $y$ is surjective for all $d\geq D$, whence ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)$ is monotonically decreasing for $d\geq D-1$. By this together with [10, Theorem 3.3.4], the homogeneous ideal $\langle F^{h}\rangle$ is zero-dimensional or trivial, i.e., there are at most a finite number of projective zeros of $F^{h}$ (and thus there are at most a finite number of affine zeros of $F$).
Remark 3.1.3.
We note that, for each $d\geq D$, the condition $\dim_{K}H_{0}(K_{\bullet})_{d-1}\!=\!\dim_{K}H_{0}(K_{\bullet})_{d}$ is equivalent to that the multiplication map $H_{0}(K_{\bullet})_{d-1}\to H_{0}(K_{\bullet})_{d}$ by $y$ is injective (and thus bijective). By this together with Lemma 2.2.2, letting $d_{0}$ be the smallest number with $d_{0}\geq D$ such that $\dim_{K}H_{0}(K_{\bullet})_{d_{0}}=\dim_{K}H_{0}(K_{\bullet})_{d_{0}+1}$, the maximal Gröbner basis degree of $\langle F^{h}\rangle$ is upper-bounded by $d_{0}$.
@@
see [25, Lemmas 5.2.1 and 5.2.2]. Note that ${\rm HF}_{R^{\prime}/(\langle F^{h}\rangle+\langle y\rangle)}(d)={\rm HF}_{R/\langle F^{\rm top}\rangle}(d)$ for any positive integer $d$. On the other hand, for $d<D$, any degree-fall does not occur, that is, if $yf\in\langle F^{h}\rangle_{d}$ with $f\in R^{\prime}$ then $f\in\langle F^{h}\rangle_{d-1}$. This can be shown by some semantic argument (see Remark 4.1.3) or also rigidly by the injectiveness of the multiplication map of $y$ in (3.1.3). Thus, we also have $\langle f\in R[y]:fy\in\langle F^{h}\rangle\rangle_{d-1}=\langle F^{h}\rangle_{d-1}$, so that
@@
Corollary 3.1.5 (Theorem 1 (3); [30, Corollary 1]).
@@
Therefore, by Theorem 2.1.4 ([14, Theorem 1]), the sequence $\bm{F}^{h}$ is $D$-regular. Here, we note that $D=\deg({\rm HS}_{R/\langle F^{\rm top}\rangle}{)}+1={\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}\right]\right)}+1$.
@@ 3.2 Solving degree for homogenized sequences
Here we assume that $\bm{F}^{\rm top}$ is semi-regular and that all degrees $d_{i}=\deg(f_{i})$ are smaller or equal to the degree of regularity $d_{\rm reg}(\langle F^{\rm top}\rangle)$. Then, any $n$-subsequence of $\bm{F}^{\rm top}$ is regular. Under this assumption, we can give a detailed discussion on the solving degree of $F^{h}$. From now on, we assume that $m\geq n$, and set $\bm{F}_{k}:=(f_{1},\ldots,f_{n+k})$ and $D_{k}:=d_{\rm reg}(\langle F_{k}^{\rm top}\rangle)$ for each $k\geq 0$. As $\bm{F}_{0}^{\rm top}$ is regular and $\bm{F}_{1}^{\rm top}$ is semi-regular, we have $D_{0}=d_{1}+\cdots+d_{n}-n+1$ and $D_{1}=\lfloor\frac{d_{1}+\cdots+d_{n+1}-n-1}{2}\rfloor+1$, see [4, Theorem 4.1]. Thus, by setting $d_{1}\leq d_{2}\leq\cdots\leq d_{m}$, we can minimize the values $D_{0}$ and $D_{1}$.
@@
Finally in this subsection, under the assumption that $\bm{F}^{\rm top}$ is cryptographic semi-regular, we show that the generalized degree of regularity of $\langle F^{h}\rangle$ (and thus solving degree of $F^{h}$) can be bounded by $D$ plus the saturation exponent, say $S_{0}$ here, that is, the minimal integer $k$ such that $\langle F\rangle^{h}=(\langle F^{h}\rangle:y^{\infty})=(\langle F^{h}\rangle:y^{k})$. See [25, p. 81] for the definition of saturation exponent.
@@
The generalized degree of regularity $\langle F^{h}\rangle$ and the solving degree of $F^{h}$ are both bounded by $D+S_{0}$.
@@
First, we show ${\rm HF}_{R^{\prime}/(\langle F^{h}\rangle+\langle y^{S_{0}}\rangle)}(d)=0$ for $d\geq D+S_{0}$, by which we have ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)={\rm HF}_{R^{\prime}/\langle F\rangle^{h}}(d-S_{0})$. Suppose for a contradiction that $(R^{\prime}/(\langle F^{h}\rangle+\langle y^{S_{0}}\rangle))_{d}\neq 0$. Then, it follows from Macaulay’s basis theorem (cf. [28, Theorem 1.5.7]) that
@@ 4 Behaviors of Gröbner bases computation
Here we show certain correspondences in the Gröbner basis computations among inputs $F^{h}$, $F^{\rm top}$, and $F$. First we revisit the correspondence among the computation of the Gröbner basis of $F^{h}$ and that of $F^{\rm top}$ given in [30, Section 5.1]. Then, we explicitly give an important correspondence between the computation of the Gröbner basis of $F^{h}$ and that of $F$, which brings an upper-bound (Lemma 4.2.4 below) on the solving degree of $F$ related to Samaev-Tenti’s bound [40].
Here we use the same notation as in the previous section, and unless otherwise noted, assume that $\bm{F}$ is cryptographic semi-regular. Let $G$, $G_{\rm hom}$, and $G_{\rm top}$ be the reduced Gröbner bases of $\langle F\rangle$, $\langle F^{h}\rangle$, and $\langle F^{\rm top}\rangle$, respectively, where their monomial orderings are DRL $\prec$ or its extension $\prec^{h}$. Also we let $D=d_{\rm reg}(\langle F^{\rm top}\rangle)$, and assume $D<\infty$. Moreover, we extend the notion of top part to a homogeneous polynomial $h$ in $R^{\prime}=R[y]$ as follows. We call $h|_{y=0}$ the top part of $h$ and denote it by $h^{\rm top}$. Thus, if $h^{\rm top}$ is not zero, it coincides with the top part $(h|_{y=1})^{\rm top}$ of the dehomogenization $h|_{y=1}$ of $h$. We remark that $g^{\rm top}=(g^{h})^{\rm top}$ for a polynomial $g$ in $R$.

### v3

@@ 2.3 Extension of the notion of degree of regularity
Note that $\prec^{\prime}$ is the homogenization $\prec^{h}$ of $\prec$. We set $f_{j}:=(h_{j})|_{y=1}$ for each $j$ with $1\leq j\leq m$, and put $F:=H|_{y=1}=\{f_{1},\ldots,f_{m}\}$ and $\bm{F}:=\bm{H}|_{y=1}=(f_{1},\ldots,f_{m})$, so that $H=F^{h}$, $\bm{H}=\bm{F}^{h}$, $H^{\rm top}=F^{\rm top}$, and $\bm{H}^{\rm top}=\bm{F}^{\rm top}$ by our assumption $(H|_{y=1})^{h}=H$. With these notations, the assertion to be proved is that we have $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\geq d_{\rm reg}(\langle F^{\rm top}\rangle_{R})$ and $\mathrm{max.GB.deg}_{\prec}(F^{\rm top})=d_{\rm reg}(\langle F^{\rm top}\rangle_{R})$. In the following, let $G_{\rm hom}$ (resp. $G_{\rm top}$) denote the reduced Gröbner basis of $\langle F^{h}\rangle_{R^{\prime}}$ (resp. $\langle F^{\rm top}\rangle_{R}$) with respect to $\prec^{h}$ (resp. $\prec$).
Put $d:=\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ and assume for a contradiction that $d<D:=d_{\rm reg}(\langle F^{\rm top}\rangle)$. By Lemma 4.1.4 below, we have ${\rm LM}(G_{\rm hom})={\rm LM}(G_{\rm top})_{\leq d}\subset R$. Also, $\langle{\rm LM}(G_{\rm hom})\rangle_{R}$ contains ${\rm LM}(G_{\rm top})$. This can be shown as follows: For any $t\in{\rm LM}(G_{\rm top})$, there are polynomials $a_{1},\ldots,a_{m}$ in $R$ such that $t={\rm LM}(\sum_{i=1}^{m}a_{i}f_{i}^{\rm top})$. In this case, it can be easily checked that $t={\rm LM}(\sum_{i=1}^{m}a_{i}f_{i}^{h})$, and thus $t$ is divisible by some element of ${\rm LM}(G_{\rm hom})$.
Thus, as $R/\langle F^{\rm top}\rangle=R/\langle G_{\rm top}\rangle$ is Artinian, it follows from $\langle{\rm LM}(G_{\rm hom})\rangle_{R}\supset{\rm LM}(G_{\rm top})$ that there is an element $g$ in $G_{\rm hom}$ with ${\rm LM}(g)=x_{n}^{d^{\prime}}$ for some $d^{\prime}\leq d$. Then, for any monomial $t\in R_{d^{\prime}}$, as $t\succeq x_{n}^{d^{\prime}}$, it belongs to $\langle{\rm LM}(\langle F^{h}\rangle)\rangle_{R^{\prime}}(=\langle{\rm LM}(G_{\rm hom})\rangle_{R^{\prime}}=\langle{\rm LM}(G_{\rm top})_{\leq d}\rangle_{R^{\prime}})$ by its weakly reverse lexicographicness. This implies that $t$ also belongs to $\langle{\rm LM}(G_{\rm top})\rangle_{R}$, and hence $R_{d^{\prime}}/\langle F^{\rm top}\rangle_{d^{\prime}}=0$. Thus, we have $d\geq d^{\prime}\geq d_{\rm reg}(\langle F^{\rm top}\rangle)$, a contradiction.
By the same argument as above, we can show that $\mathrm{max.GB.deg}_{\prec}(F^{\rm top})=d_{\rm reg}(\langle F^{\rm top}\rangle)$ as follows. Recall from Lemma 2.1.8 that $\mathrm{max.GB.deg}_{\prec}(F^{\rm top})\leq d_{\rm reg}(\langle F^{\rm top}\rangle)$. Thus, we assume to the contrary that $d:=\mathrm{max.GB.deg}_{\prec}(F^{\rm top})<D:=d_{\rm reg}(\langle F^{\rm top}\rangle)$. Then, it follows from Lemma 4.1.4 below that ${\rm LM}(G_{\rm top})={\rm LM}(G_{\rm hom})_{\leq d}$ and the ideal $\langle{\rm LM}(G_{\rm top})\rangle_{R}$ has the weak reverse lexicographicness up to $d$. Since there is an element $g$ in $G_{\rm top}$ with ${\rm LM}(g)=x_{n}^{d^{\prime}}$ for some $d^{\prime}\leq d$, any monomial in $R_{d^{\prime}}$ belongs to $\langle{\rm LM}(G_{\rm top})\rangle$ and so $R_{d^{\prime}}/\langle F^{\rm top}\rangle_{d^{\prime}}=0$, which implies $d\geq d^{\prime}\geq D$, a contradiction. ∎
@@
In Lemma 2.3.8, when $R/\langle H^{\rm top}\rangle$ is Artinian, we can easily prove the equality $\mathrm{max.GB.deg}_{\prec}(H^{\rm top})=d_{\rm reg}(\langle H^{\rm top}\rangle)$ if $\langle\mathrm{LM}(\langle H^{\rm top}\rangle)\rangle$ is weakly reverse lexicographic, not assuming that $\bm{H}^{\rm top}$ is cryptographic semi-regular nor that $\langle\mathrm{LM}(\langle H\rangle)\rangle$ is weakly reverse lexicographic.
@@ 3.1 Bounded regularity of homogenized sequences
Here we revisit the main results in [31, Section 4]. For the readability, we remain the proofs. Also, as additional remarks, we explicitly give an important property of the Hilbert-Poincaré series of $R^{\prime}/\langle F^{h}\rangle$ with $R^{\prime}=R[y]$, and also give an alternative proof for [31, Theoem 7] (Theorem 3.1.1 below).
@@
Theorem 3.1.1 (Theorem 1 (1); [31, Theoem 7]).
@@
${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)={\rm HF}_{R/\langle F^{\rm top}\rangle}(d)+{\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d-1),$ (3.1.1)
@@
$K_{i}\otimes_{R^{\prime}}R\cong\bigoplus_{1\leq j_{1}<\cdots<j_{i}\leq m}R(-d_{j_{1}\cdots j_{i}})\mathbf{e}_{j_{1}\cdots j_{i}},$
@@
by the dimension theorem. Since $H_{0}(K_{\bullet})=R^{\prime}/\langle F^{h}\rangle$ and $H_{0}(K_{\bullet}\otimes_{R^{\prime}}R)=R/\langle F^{\rm top}\rangle$, we have the equality (3.1.1), as desired. ∎
Remark 3.1.2 (Theorem 1 (2), (3); [31, Remark 6]).
Note that, in the proof of Theorem 3.1.1, the multiplication map $H_{0}(K_{\bullet})_{d-1}\to H_{0}(K_{\bullet})_{d}$ by $y$ is injective for all $d<D$, whence ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)$ is monotonically increasing for $d<D$. On the other hand, since $H_{0}(K_{\bullet}\otimes_{R^{\prime}}R)_{d}=(R/\langle F^{\rm top}\rangle)_{d}=0$ for all $d\geq D$ by the definition of the degree of regularity, the multiplication map $H_{0}(K_{\bullet})_{d-1}\to H_{0}(K_{\bullet})_{d}$ by $y$ is surjective for all $d\geq D$, whence ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)$ is monotonically decreasing for $d\geq D-1$. By this together with [11, Theorem 3.3.4], the homogeneous ideal $\langle F^{h}\rangle$ is zero-dimensional or trivial, i.e., there are at most a finite number of projective zeros of $F^{h}$ (and thus there are at most a finite number of affine zeros of $F$).
@@
see [26, Lemmas 5.2.1 and 5.2.2]. Note that ${\rm HF}_{R^{\prime}/(\langle F^{h}\rangle+\langle y\rangle)}(d)={\rm HF}_{R/\langle F^{\rm top}\rangle}(d)$ for any positive integer $d$. On the other hand, for $d<D$, any degree-fall does not occur, that is, if $yf\in\langle F^{h}\rangle_{d}$ with $f\in R^{\prime}$ then $f\in\langle F^{h}\rangle_{d-1}$. This can be shown by some semantic argument (see Remark 4.1.3) or also rigidly by the injectiveness of the multiplication map of $y$ in (3.1.3). Thus, we also have $\langle f\in R[y]:fy\in\langle F^{h}\rangle\rangle_{d-1}=\langle F^{h}\rangle_{d-1}$, so that
@@
Corollary 3.1.4 (Theorem 1 (3); [31, Corollary 1]).
@@
Therefore, by Theorem 2.1.4 ([15, Theorem 1]), the sequence $\bm{F}^{h}$ is $D$-regular. Here, we note that $D=\deg({\rm HS}_{R/\langle F^{\rm top}\rangle}{)}+1={\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}\right]\right)}+1$.
@@ 3.2 Solving degree for homogenized sequences
Here we assume that $\bm{F}^{\rm top}$ is semi-regular and that all degrees $d_{i}=\deg(f_{i})$ are smaller or equal to the degree of regularity $d_{\rm reg}(\langle F^{\rm top}\rangle)$. Then, any $n$-subsequence of $\bm{F}^{\rm top}$ is regular. Under this assumption, we can give a detailed discussion on the solving degree of $F^{h}$. From now on, we assume that $m\geq n$, and set $\bm{F}_{k}:=(f_{1},\ldots,f_{n+k})$ and $D_{k}:=d_{\rm reg}(\langle F_{k}^{\rm top}\rangle)$ for each $k\geq 0$. As $\bm{F}_{0}^{\rm top}$ is regular and $\bm{F}_{1}^{\rm top}$ is semi-regular, we have $D_{0}=d_{1}+\cdots+d_{n}-n+1$ and $D_{1}=\lfloor\frac{d_{1}+\cdots+d_{n+1}-n-1}{2}\rfloor+1$, see [5, Theorem 4.1]. Thus, by setting $d_{1}\leq d_{2}\leq\cdots\leq d_{m}$, we can minimize the values $D_{0}$ and $D_{1}$.
@@
Finally in this subsection, under the assumption that $\bm{F}^{\rm top}$ is cryptographic semi-regular, we show that the generalized degree of regularity of $\langle F^{h}\rangle$ (and thus solving degree of $F^{h}$) can be bounded by $D$ plus the saturation exponent, say $S_{0}$ here, that is, the minimal integer $k$ such that $\langle F\rangle^{h}=(\langle F^{h}\rangle:y^{\infty})=(\langle F^{h}\rangle:y^{k})$. See [26, p. 81] for the definition of saturation exponent.
@@
The generalized degree of regularity of $\langle F^{h}\rangle$ and the solving degree of $F^{h}$ are both bounded by $D+S_{0}$.
@@
First, we show ${\rm HF}_{R^{\prime}/(\langle F^{h}\rangle+\langle y^{S_{0}}\rangle)}(d)=0$ for $d\geq D+S_{0}$, by which we have ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)={\rm HF}_{R^{\prime}/\langle F\rangle^{h}}(d-S_{0})$. Suppose for a contradiction that $(R^{\prime}/(\langle F^{h}\rangle+\langle y^{S_{0}}\rangle))_{d}\neq 0$. Then, it follows from Macaulay’s basis theorem (cf. [29, Theorem 1.5.7]) that
@@ 4 Behaviors of Gröbner bases computation
We use the same notation as in Section 3. Here we show certain correspondences in the Gröbner basis computations among inputs $F^{h}$, $F^{\rm top}$, and $F$. First we revisit the correspondence among the computation of the Gröbner basis of $F^{h}$ and that of $F^{\rm top}$ given in [31, Section 5.1]. Then, we explicitly give an important correspondence between the computation of the Gröbner basis of $F^{h}$ and that of $F$, which brings an upper-bound (Lemma 4.2.4 below) on the solving degree of $F$ related to Samaev-Tenti’s bound [41]. Subsequently, we observe actual behavior of Gröbner basis computation from a part of our experimental results. From our observation, we estimate the complexity of the Gröbner basis computation for $F^{h}$ (and $F$), and will raise a conjecture and some questions related to the generalized cryptographic semi-regularity of $\bm{F}^{h}$.
Here we use the same notation as in the previous section, and unless otherwise noted, assume that $\bm{F}$ is cryptographic semi-regular. Let $G$, $G_{\rm hom}$, and $G_{\rm top}$ be the reduced Gröbner bases of $\langle F\rangle$, $\langle F^{h}\rangle$, and $\langle F^{\rm top}\rangle$, respectively, where their monomial orderings are DRL $\prec$ or its homogenization $\prec^{h}$. Also we let $D=d_{\rm reg}(\langle F^{\rm top}\rangle)$, and assume $D<\infty$. Moreover, we use the notion of top part to a homogeneous polynomial $h$ in $R^{\prime}=R[y]$, see Definition 2.3.4.